# Coupling Theory, Optimal Transport, and Strassen’s Theorem Beyond Regular Orders

### v1

@@ Abstract.
coupling; duality; equivalence relations; Kantorovich duality; optimal transport; partial orders; stochastic domination; Strassen’s theorem
@@ 1.1. Background
A coupling of Borel probability measures $\mu,\nu$ on a Polish space $S$ is a Borel probability measure on $S\times S$ satisfying $\pi(S\times\,\cdot\,)=\mu$ and $\pi(\,\cdot\,\times S)=\nu$. Equivalently, a coupling is the joint law of a pair of $S$-valued random variables $(X,Y)$ defined on a common probability space in such a way that $X$ has distribution $\mu$ and $Y$ has distribution $\nu$. Coupling is one of the core techniques in probability theory, where it is used ubiquitously in essentially all sub-disciplines; see [11, 43, 33] for accessible introductions. At the same time, the study of couplings is interesting in its own right and a great deal literature has addressed fundamental questions in coupling theory, particularly in the setting of stochastic processes [19, 34, 18, 41, 13, 21, 38, 8, 30, 25, 9, 5, 26, 27, 7, 10, 4, 29, 22] but also more generally [17, 39, 23].
@@ 1.2. Problem Statement
Towards making a rigorous proof out of this formal argument, we consider the setting above and we apply the result [35, Corollary 2.3.9] to get
@@
Returning to the present work, our goal is investigate Thorisson’s working hypothesis for reflexive transitive relations that need not be closed sets, but can be written as a countable increasing union of closed sets; that is, we focus on reflexive, transitive, $F_{\sigma}$ relations. One motivation for this setting is the prior literature [23, 36] studying an equivalence relation $\cong$, which show that strong duality holds under some purely measure-theoretic conditions on $\cong$. Notably, these works show that strong duality holds in some settings where standard approaches based on Kantorovich duality [37] can fail. Another motivation for our setting concerns applications of coupling theory in probability. Indeed, many interesting coupling problems involve equivalence relations $\cong$ such that $R_{\cong}$ is not closed (see [23, Section 2] for several examples), and in Subsection 1.4 we discuss the coupling problem of eventual domination which involves a partial order $\preceq$ such that $R_{\preceq}$ is not closed. As such, it is natural to investigate Strassen’s theorem for $F_{\sigma}$ partial orders. The simplest setting capturing both of these motivations is reflexive, transitive, $F_{\sigma}$ relations.
@@ 1.3. Main Results
The proof of Theorem 1.1, about which we now make a few remarks, is given in Section 2. First, we note that the negative part is shown by example, namely in Subsection 2.4 we construct a reflexive transitive $F_{\sigma}$ relation on $[0,2]$ and two Borel probability measures $\mu,\nu$ on $[0,2]$ such that (1.1) has value zero but admits no solutions. Second, we note that the positive part involves a few steps, one of which is a general result in optimal transport that may be of independent interest. To state it, recall that the usual optimal transport theory (e.g., [45]) requires a cost function $c:S\times S\to\mathbb{R}$ which is lower semicontinuous. The following result shows that the Kantorovich dual problem is still well-behaved when $c$ lacks this regularity.
@@ 1.4. Future Work
We now describe some open questions related to this work.
First, there are questions arising from the connection we have identified between Thorisson’s working hypothesis and optimal transport for zero-one-valued cost functions. For instance, one may hope to understand cyclic monotonicity for such cost functions, hence [6, Theorem 1] establishing optimality conditions for the resulting optimal transport problem; this would imply some further structure of the various couplings statements appearing as statement (ii) above, as was previously done for equivalence relations in [23, Proposition 3.12]. For another example, one may hope to understand the existence of Monge solutions to the resulting optimal transport problem; this would imply that the couplings appearing in statement (ii) above can be achieved “without additional randomness,” as was previously done in [2] in the setting of measurable selections of random set-valued functions and in [22] in the setting of germ couplings of Brownian motions with drift.
Second, there are questions concerning applications of our general results to some problems of interest in probability theory. For instance, one may consider the problem of characterizing when two real-valued stochastic process $\{X_{t}\}_{t\geq 0}$ and $\{Y_{t}\}_{t\geq 0}$ can be coupled to satisfy eventual domination, meaning that almost surely we have $X_{t}\leq Y_{t}$ for sufficiently large (possibly random) $t\geq 0$. Eventual domination may of course be achieved by a coupling satisfying everywhere domination (as in Strassen’s theorem) or eventual equality (as in the classical works [19, 34, 18]), but such sufficient conditions are likely too weak to yield complete characterizations. We believe this problem is particularly interesting when $\{X_{t}\}_{t\geq 0}$ and $\{Y_{t}\}_{t\geq 0}$ are transient Markov processes, in which case it is natural to conjecture that eventual domination should be closely related to stochastic domination of the induced laws on the common Martin boundary of the two processes. More concretely, one may aim to show that the existence of a coupling satisfying eventual domination for, say, two Lévy processes, is equivalent to some explicit property of their characteristic triplets or their Laplace exponents.
@@ 2.1. Dual variable elimination
Note that $R$ being $F_{\sigma}$ implies that there exists a non-decreasing sequence $\{R_{n}\}_{n\in\mathbb{N}}$ of closed sets in $S\times S$ such that $R=\bigcup_{n\in\mathbb{N}}R_{n}$. Then, let us write $c=1-\mathbf{1}_{R}$ and $c_{n}=1-\mathbf{1}_{R_{n}}$ for each $n\in\mathbb{N}$. It follows that the hypotheses of Theorem 1.2 are satisfied, so, by combining with general results on Kantorovich duality (e.g., [35, Corollary 2.3.9]), we have
@@
Let $\phi,\psi:S\to\mathbb{R}$ be arbitrary bounded, Borel measurable functions satisfying $\phi(x)+\psi(y)\leq c(x,y)$ for all $x,y\in S$, and let us write
$\psi^{c}(x)\coloneqq\inf_{y\in S}(c(x,y)-\psi(y)),$
for the $c$-transform of $\psi$. Now note that property (a) implies $\psi^{c}\leq-\psi$ pointwise, and that property (b) implies $\psi^{cc}=-\psi^{c}$. Moreover, by standard optimal transport considerations, we have
@@
As projections of Borel sets, these sub-level sets are universally measurable. (See [24, Theorem A1.2].)
@@
$\int g\,\,\textnormal{d}\mu$ $=\sum_{i=1}^{r}b_{i}\mu(K_{i})$
$\geq\sum_{i=1}^{r}b_{i}\left(\mu(E_{i})-\frac{\varepsilon}{2rb_{i}}\right)$
$=\int\eta\,\,\textnormal{d}\mu-\frac{\varepsilon}{2}$
$\geq\int\psi^{c}\,\,\textnormal{d}\mu-\varepsilon.$
@@
$\geq\int\psi^{c}\,\,\textnormal{d}\mu-\int\psi^{c}\,\,\textnormal{d}\nu-\varepsilon$
$\geq\int\phi\,\,\textnormal{d}\mu-\int\psi\,\,\textnormal{d}\mu-\varepsilon$
@@
For $y$ such that $g(y)=0$, i.e. $y\notin\bigcup_{i=1}^{r}K_{i}$, the condition becomes $c_{n}(x,y)<\alpha\leq 0$, which is impossible because $c_{n}\geq 0$. So we may assume that $y\in K_{1}\cup\cdots\cup K_{r}$.
$=p\left(\bigcup_{n=1}^{\infty}\bigcup_{i=1}^{r}\{(x,y)\in S\times K_{i}:c_{n}(x,y)-b_{i}<\alpha\}\right)$
$=\bigcup_{n=1}^{\infty}\bigcup_{i=1}^{r}p(\{(x,y)\in S\times K_{i}:c_{n}(x,y)<\alpha+b_{i}\})$
@@
Discussions about the Borel measurability of $c$-transformed functions appear already in most standard introductions to optimal transport (e.g., [45, Remark 5.5 and p. 69]). Our result required slightly more work, since our cost function $c$ need not even be lower semi-continuous.
@@ 2.2. Existence of dual maximizer
The claim (b) implies (a) is immediate, so we only need to prove that (a) implies (b). Let $E:=\{(x,y)\in S\times S:f(x)+g(y)>\ell(x,y)\}$, and note by assumption that we have $\pi(E)=0$ for all $\pi\in\Pi(\mu,\nu)$. By [20, Corollary, p. 500] this implies that there exist Borel sets $N_{f},N_{g}\subseteq S$ with $\mu(N_{f})=\nu(N_{g})=0$ such that $E\subseteq(N_{f}\times S)\cup(S\times N_{g})$. Now we set
@@
$h(x,y):=1-\mathbf{1}_{R}(x,y)-\phi_{\infty}(x)+\phi_{\infty}(y),$
as well as
@@
Lemma 2.3 now shows that we can adjust $\phi_{\infty}$$\phi_{\infty}$ by setting them to be 0 on $\rho$-null sets $N_{\phi},N_{-\phi}$ to obtain Borel-measurable versions $\phi_{\infty}^{\prime},(-\phi_{\infty})^{\prime}$ which satisfy the cost inequality pointwise; by making the modification using $N_{\phi}\cup N_{-\phi}$ instead for both functions, we can guarantee that $(-\phi_{\infty})^{\prime}=-\phi_{\infty}^{\prime}$ pointwise. And since the set $N_{\phi}\cup N_{-\phi}$ is $\rho$-null and hence both $\mu$-null and $\nu$-null (this is where we use the choice of $L^{2}(\rho)\times L^{2}(\rho)$ rather than $L^{2}(\mu)\times L^{2}(\nu)$),
@@
$=\lim_{k\to\infty}\int\phi_{n_{k}}\,\,\textnormal{d}\mu-\int\phi_{n_{k}}\,\,\textnormal{d}\nu$
@@
$=\sup_{\begin{subarray}{c}\phi\in b\mathcal{B}(S)\\ \phi\ominus\phi\leq c\end{subarray}}\int\phi\,\,\textnormal{d}\mu-\int\phi\,\,\textnormal{d}\nu,$
where the third inequality follows from weak convergence, paired with the fact that $\frac{\,\textnormal{d}\mu}{\,\textnormal{d}\rho}-\frac{\,\textnormal{d}\nu}{\,\textnormal{d}\rho}\in L^{2}(\rho)$ (in fact, $0\leq\frac{\,\textnormal{d}\mu}{\,\textnormal{d}\rho},\frac{\,\textnormal{d}\nu}{\,\textnormal{d}\rho}\leq 2$).
@@ 2.3. Conversion to indicator maximizer
where we have defined
$\mathbf{1}_{A_{t}}:=\{x\in S:t\leq\phi_{\infty}^{\prime}(x)\leq 1\}$
for all $0\leq t\leq 1$.
@@ 2.4. Non-existence of primal minimizers
The only remaining step of the proof is to show that the primal problem, i.e. the right side of (2.1), can fail to admit minimizers. Towards this end, we give the following elementary construction.
@@
First, we claim that the primal value is zero. To see this, fix $0<\varepsilon<1$ and define the map $T_{\varepsilon}(x):[0,1]\to[1,2]$ via
$T(x):=1+[x-\varepsilon],$
where $\alpha]$ denotes the fractional part of a real number $\alpha\in\mathbb{R}$; then define $\pi_{\varepsilon}\in\Pi(\mu,\nu)$ as the pushforward $\pi_{\varepsilon}:=(\textnormal{id},T_{\varepsilon})_{\#}\mu$. Note that $\pi_{\varepsilon}(R)=1-\varepsilon$ for all $0<\varepsilon<1$, so taking $\varepsilon\to 0$ shows indeed that the primal value is zero.
@@
Example 2.4 is counterexample to the claim “Thorisson’s working hypothesis holds for $F_{\sigma}$ partial orders,” although only insofar as the primal problem may fail to admit minimizers. In light of analogous work on equivalence relations [23, Proposition 3.18], it is natural to wonder whether the weaker claim “Thorisson’s working hypothesis holds for partial orders that can be written as a countable increasing union of closed partial orders” is true; note, however, that Example 2.4 is a counterexample to this statement as well, since the closed sets in (2.3) are indeed partial orders. This discrepancy highlights that there is further structure in equivalence relations, compared to in partial orders, which allows for a more complete duality theory.
Although the partial order $([0,2],R)$ from Example 2.4 may seem artificial, we note that it can be embedded into many other partial orders of interest, hence establishing counterexamples in other settings. Of particular interest is the partial order of eventual domination from Subsection 1.4, which we denote $(\mathbb{R}^{\mathbb{N}},\leq^{\ast})$; we are not aware of how to explicitly construct such an embedding from $([0,2],R)$ to $(\mathbb{R}^{\mathbb{N}},\leq^{\ast})$, but we believe it may be achievable with further work. However, it appears that the embedability of certain partial orders into $(\mathbb{R}^{\mathbb{N}},\leq^{\ast})$ is a delicate set-theoretic question and may be independent of ZFC (e.g., [14, Lemma 4.1]).

### v2

@@ Abstract.
coupling; duality; equivalence relations; eventual domination; Kantorovich duality; optimal transport; partial orders; stochastic domination; Strassen’s theorem
@@ 1.1. Background
A coupling of Borel probability measures $\mu,\nu$ on a Polish space $S$ is a Borel probability measure on $S\times S$ satisfying $\pi(S\times\,\cdot\,)=\mu$ and $\pi(\,\cdot\,\times S)=\nu$. Equivalently, a coupling is the joint law of a pair of $S$-valued random variables $(X,Y)$ defined on a common probability space in such a way that distributions of $X$,$Y$ are $\mu$$\nu$, respectively. Coupling is one of the core techniques in probability theory, where it is used ubiquitously in essentially all sub-disciplines; see [7, 33, 23] for introductions. At the same time, the study of couplings is interesting in its own right and a great deal literature has addressed fundamental questions in coupling theory, particularly in the setting of stochastic processes [13, 24, 12, 31, 9, 15, 28, 6, 5, 3, 20, 16] but also more generally [11, 29, 17].
@@ 1.2. Problem Statement
Towards making a rigorous proof out of this formal argument, we consider the setting above and we apply the result [25, Corollary 2.3.9] to get
@@
Some important relations (from the point of view of probability theory) do not correspond to closed sets and yet similar coupling results are known. For example, the equivalence relation of eventual equality or tail equivalence of sequences or paths has been studied in [13, 24, 12] and is closely related to some aspects of probabilistic potential theory and Markov chain mixing times. Another example is the equivalence relation of germ equivalence which has been studied in [9, 17]. A motivating example for the present work, which has not been previously studied, is the relation of eventual domination or tail domination, say on $S=\mathbb{R}^{\mathbb{N}}$ or more generally on the space of sequences on any partially ordered set; concretely, let $\leq^{\ast}$ be the relation on $S$ whereby $\{x_{n}\}_{n\in\mathbb{N}}\leq^{\ast}\{y_{n}\}_{n\in\mathbb{N}}$ if and only if $x_{n}\leq y_{n}$ for all sufficiently large $n\in\mathbb{N}$. In terms of probability theory, a coupling result for eventual domination à la Thorisson would allow one to characterize the possibility of long-term comparisons of, say, a Markov chain started from different initializations or of different Markov processes started from the same initialization.
@@
Returning to the present work, our goal is investigate Thorisson’s working hypothesis for reflexive, transitive relations that need not be closed sets, but can be written as a countable increasing union of closed sets; that is, we focus on reflexive, transitive, $F_{\sigma}$ relations. As we have discussed, we have three motivations for this level of generality. First is the prior literature [17, 26] studying an equivalence relation $\cong$, which show that strong duality holds under some purely measure-theoretic conditions on $\cong$; notably, these works show that strong duality holds in some settings where standard approaches based on Kantorovich duality [27] can fail. Second is the natural question of whether Strassen’s theorem for closed partial order extends to $F_{\sigma}$ partial orders. Third is providing new results for the problem of eventual domination, which is reflexive and transitive but not antisymmetric (unless passing to a suitable quotient).
@@ 1.3. Main Results
The proof of Theorem 1.1, about which we now make a few remarks, is given in Section 2. First, we note that the negative part is shown, in Subsection 2.4, by directly constructing counterexamples in the setting of a simple $F_{\sigma}$ partial order on $[0,2]$ (Example 2.4) and the pre-order of eventual domination on $\mathbb{R}^{\mathbb{N}}$ (Example 2.5). Second, we note that the positive part involves a few steps, one of which is a general result in optimal transport that may be of independent interest. To state it, recall that the usual optimal transport theory (e.g., [34]) requires a cost function $c:S\times S\to\mathbb{R}$ which is lower semicontinuous. The following result shows that the Kantorovich dual problem is still well-behaved when $c$ lacks this regularity.
@@ 1.4. Future Work
First, there are questions arising from the connection we have identified between Thorisson’s working hypothesis and optimal transport for zero-one-valued cost functions. For instance, one may hope to understand cyclic monotonicity for such cost functions, hence [4, Theorem 1] establishing optimality conditions for the resulting optimal transport problem; this would imply some further structure of the various couplings statements appearing as statement (ii) above, as was previously done for equivalence relations in [17, Proposition 3.12]. For another example, one may hope to understand the existence of Monge solutions to the resulting optimal transport problem; this would imply that the couplings appearing in statement (ii) above can be achieved “without additional randomness,” as was previously done in [1] in the setting of measurable selections of random set-valued functions and in [16] in the setting of germ couplings of Brownian motions with drift.
Second, there are questions concerning applications of our general results to some problems of interest in probability theory. For instance, one may consider the problem of characterizing when two real-valued stochastic process $\{X_{t}\}_{t\geq 0}$ and $\{Y_{t}\}_{t\geq 0}$ can be coupled to satisfy eventual domination. We believe this problem is particularly interesting when $\{X_{t}\}_{t\geq 0}$ and $\{Y_{t}\}_{t\geq 0}$ are transient Markov processes, where a natural conjecture is that eventual domination should be closely related to stochastic domination of the induced laws on the common Martin boundary of the two processes. More concretely, one may aim to show that the existence of a coupling satisfying eventual domination for, say, two Lévy processes, is equivalent to some explicit property of their characteristic triplets or their Laplace exponents.
@@ 2.1. Dual variable elimination
Note that $R$ being $F_{\sigma}$ implies that there exists a non-decreasing sequence $\{R_{n}\}_{n\in\mathbb{N}}$ of closed sets in $S\times S$ such that $R=\bigcup_{n\in\mathbb{N}}R_{n}$. Then, let us write $c=1-\mathbf{1}_{R}$ and $c_{n}=1-\mathbf{1}_{R_{n}}$ for each $n\in\mathbb{N}$. It follows that the hypotheses of Theorem 1.2 are satisfied, so, by combining with general results on Kantorovich duality (e.g., [25, Corollary 2.3.9]), we have
@@
Let $\phi,\psi:S\to\mathbb{R}$ be arbitrary bounded, Borel measurable functions satisfying $\phi(x)+\psi(y)\leq c(x,y)$ for all $x,y\in S$, and write $\psi^{c}(x)\coloneqq\inf_{y\in S}(c(x,y)-\psi(y))$ for the $c$-transform of $\psi$. Now note that property (a) implies $\psi^{c}\leq-\psi$ pointwise, and that property (b) implies $\psi^{cc}=-\psi^{c}$. Moreover, by standard optimal transport considerations, we have
@@
As projections of Borel sets, these sub-level sets are universally measurable. (See [18, Theorem A1.2].)
@@
$\int g\,\,\textnormal{d}\mu=\sum_{i=1}^{r}b_{i}\mu(K_{i})\geq\sum_{i=1}^{r}b_{i}\left(\mu(E_{i})-\frac{\varepsilon}{2rb_{i}}\right)=\int\eta\,\,\textnormal{d}\mu-\frac{\varepsilon}{2}\geq\int\psi^{c}\,\,\textnormal{d}\mu-\varepsilon.$
@@
$\geq\int\psi^{c}\,\,\textnormal{d}\mu-\int\psi^{c}\,\,\textnormal{d}\nu-\varepsilon\geq\int\phi\,\,\textnormal{d}\mu-\int\psi\,\,\textnormal{d}\mu-\varepsilon$
@@
For $y$ such that $g(y)=0$, i.e. $y\notin\bigcup_{i=1}^{r}K_{i}$, the condition becomes $c_{n}(x,y)<\alpha\leq 0$, which is impossible because $c_{n}\geq 0$. So we may assume that $y\in K_{1}\cup\cdots\cup K_{r}$. Then
$(g^{c})^{-1}((-\infty,\alpha))$ $=p\left(\bigcup_{n=1}^{\infty}\bigcup_{i=1}^{r}\{(x,y)\in S\times K_{i}:c_{n}(x,y)-b_{i}<\alpha\}\right)$
$=\bigcup_{n=1}^{\infty}\bigcup_{i=1}^{r}p(\{(x,y)\in S\times K_{i}:c_{n}(x,y)<\alpha+b_{i}\}).$
@@
Discussions about the Borel measurability of $c$-transformed functions appear already in most standard introductions to optimal transport (e.g., [34, Remark 5.5 and p. 69]). Our result required slightly more work, since our cost function $c$ need not even be lower semi-continuous.
@@ 2.2. Existence of dual maximizer
The claim (b) implies (a) is immediate, so we only need to prove that (a) implies (b). Let $E:=\{(x,y)\in S\times S:f(x)+g(y)>\ell(x,y)\}$, and note by assumption that we have $\pi(E)=0$ for all $\pi\in\Pi(\mu,\nu)$. By [14, Corollary, p. 500] this implies that there exist Borel sets $N_{f},N_{g}\subseteq S$ with $\mu(N_{f})=\nu(N_{g})=0$ such that $E\subseteq(N_{f}\times S)\cup(S\times N_{g})$. Now we set
@@
$h(x,y)$ $:=1-\mathbf{1}_{R}(x,y)-\phi_{\infty}(x)+\phi_{\infty}(y),\qquad\textnormal{ and}$
@@
Lemma 2.3 now shows that we can adjust $\phi_{\infty}$$\phi_{\infty}$ by setting them to be 0 on $\rho$-null sets $N_{\phi},N_{-\phi}$ to obtain Borel-measurable versions $\phi_{\infty}^{\prime},(-\phi_{\infty})^{\prime}$ which satisfy the cost inequality pointwise; by making the modification using $N_{\phi}\cup N_{-\phi}$ instead for both functions, we can guarantee that $(-\phi_{\infty})^{\prime}=-\phi_{\infty}^{\prime}$ pointwise. And since the set $N_{\phi}\cup N_{-\phi}$ is $\rho$-null and hence both $\mu$-null and $\nu$-null (this is where we use the choice of $L^{2}(\rho)\times L^{2}(\rho)$ rather than $L^{2}(\mu)\times L^{2}(\nu)$), we can use weak convergence to get
@@
where we used $0\leq\frac{\,\textnormal{d}\mu}{\,\textnormal{d}\rho},\frac{\,\textnormal{d}\nu}{\,\textnormal{d}\rho}\leq 2$ hence $\frac{\,\textnormal{d}\mu}{\,\textnormal{d}\rho}-\frac{\,\textnormal{d}\nu}{\,\textnormal{d}\rho}\in L^{2}(\rho)$. Then
$\int\phi_{\infty}^{\prime}\,\,\textnormal{d}\mu-\int\phi_{\infty}^{\prime}\,\,\textnormal{d}\nu$ $=\lim_{k\to\infty}\int\phi_{n_{k}}\,\,\textnormal{d}\mu-\int\phi_{n_{k}}\,\,\textnormal{d}\nu$
@@
$=\sup_{\begin{subarray}{c}\phi\in b\mathcal{B}(S)\\ \phi\ominus\phi\leq c\end{subarray}}\int\phi\,\,\textnormal{d}\mu-\int\phi\,\,\textnormal{d}\nu.$
@@ 2.3. Conversion to indicator maximizer
where we have defined $\mathbf{1}_{A_{t}}:=\{x\in S:t\leq\phi_{\infty}^{\prime}(x)\leq 1\}$ for all $0\leq t\leq 1$.
@@ 2.4. Non-existence of primal minimizers
The only remaining step of the proof is to show that the primal problem, i.e. the right side of (2.1), can fail to admit minimizers. Towards this end, we construct two counterexamples; Example 2.4 is elementary, and Example 2.5 is an embedding of Example 2.4 into a more complicated setting of interest.
@@
First, we claim that the primal value is zero. To see this, fix $0<\varepsilon<1$ and define the map $T_{\varepsilon}(x):[0,1]\to[1,2]$ via $T(x):=1+[x-\varepsilon]$, where $\alpha]$ denotes the fractional part of a real number $\alpha\in\mathbb{R}$; then define $\pi_{\varepsilon}\in\Pi(\mu,\nu)$ as the pushforward $\pi_{\varepsilon}:=(\textnormal{id},T_{\varepsilon})_{\#}\mu$. Note $\pi_{\varepsilon}(R)=1-\varepsilon$ for all $0<\varepsilon<1$, so taking $\varepsilon\to 0$ shows that the primal value is zero.
@@
Example 2.4 is counterexample to the claim “Thorisson’s working hypothesis holds for $F_{\sigma}$ partial orders,” although only insofar as the primal problem may fail to admit minimizers. In light of analogous work on equivalence relations [17, Proposition 3.18], it is natural to wonder whether the weaker claim “Thorisson’s working hypothesis holds for partial orders that can be written as a countable increasing union of closed partial orders” is true; note, however, that Example 2.4 is a counterexample to this statement as well, since the closed sets in (2.3) are indeed partial orders. This discrepancy highlights that there is further structure in equivalence relations, compared to in partial orders, which allows for a more complete duality theory.
Although the partial order $([0,2],R)$ from Example 2.4 may seem artificial, we note that it can be embedded into many other orders of interest, hence establishing counterexamples in other settings. Of particular interest is the pre-order of eventual domination from Subsection 1.4, which we denote $(\mathbb{R}^{\mathbb{N}},\leq^{\ast})$.
Example 2.5.
Let $S=\mathbb{R}^{\mathbb{N}}$, let $\leq^{\ast}$ denote the pre-order of eventual domination, and let $([0,2],R)$ be the partially ordered set defined in Example 2.4. Our goal is to construct a function $\phi$ such that $\phi:([0,2],R)\to(\mathbb{R}^{\mathbb{N}},\leq^{\ast})$ is an order embedding, meaning we have $(x,y)\in R$ if and only if $\phi(x)\leq^{\ast}\phi(y)$. Indeed, we may set
$\phi(z):=\begin{cases}a(z)&\textnormal{ if }0\leq z\leq 1\\ b(z-1)&\textnormal{ if }1<z\leq 2\\ \end{cases},\qquad\textnormal{where}$
[…]
It can be checked by casework that $\phi$ is an order embedding. In order to transfer Example 2.4 to the present setting, we let $\mu$ and $\nu$ be the Lebesgue measures on the intervals $[0,1]$ and $[1,2]$, respectively, and then push these forward to $\tilde{\mu}:=\phi_{\#}\mu$ and $\tilde{\nu}:=\phi_{\#}\nu$ on $\mathbb{R}^{\mathbb{N}}$.
First, we claim that the primal problem has value zero. To see this, note that for any $\pi\in\Pi(\mu,\nu)$ we have
$\int_{S\times S}(1-\mathbf{1}_{R_{\leq^{\ast}}})\,\textnormal{d}\left((\phi,\phi)_{\#}\pi\right)=\int_{[0,2]\times[0,2]}(1-\mathbf{1}_{R})\,\textnormal{d}\pi.$
Hence, an $\varepsilon$-optimal coupling for the primal problem in $([0,2],R)$ leads to an $\varepsilon$-optimal coupling for the primal problem in $(S,\leq^{\ast})$, and we conclude that the optimal value is zero.
[…]
$\int_{[0,2]\times[0,2]}(1-\mathbf{1}_{R})\,\textnormal{d}\pi=\int_{S\times S}(1-\mathbf{1}_{R_{\leq^{\ast}}})\,\textnormal{d}\tilde{\pi}=0.$
This contradicts Example 2.4 since no such $\pi$ can exist.